# TADT-CSA: Temporal Advantage Decision Transformer with Contrastive State Abstraction for Generative Recommendation

### v1

@@ Abstract
With the rapid advancement of Transformer-based Large Language Models (LLMs), generative recommendation has shown great potential in enhancing both the accuracy and semantic understanding of modern recommender systems. Compared to LLMs, the Decision Transformer (DT) is a lightweight generative model applied to sequential recommendation tasks. However, DT faces challenges in trajectory stitching, often producing suboptimal trajectories. Moreover, due to the high dimensionality of user states and the vast state space inherent in recommendation scenarios, DT can incur significant computational costs and struggle to learn effective state representations. To overcome these issues, we propose a novel Temporal Advantage Decision Transformer with Contrastive State Abstraction (TADT-CSA) model. Specifically, we combine the conventional Return-To-Go (RTG) signal with a novel temporal advantage (TA) signal that encourages the model to capture both long-term returns and their sequential trend. Furthermore, we integrate a contrastive state abstraction module into the DT framework to learn more effective and expressive state representations. Within this module, we introduce a TA–conditioned State Vector Quantization (TAC-SVQ) strategy, where the TA score guides the state codebooks to incorporate contextual token information. Additionally, a reward prediction network and a contrastive transition prediction (CTP) network are employed to ensure the state codebook preserves both the reward information of the current state and the transition information between adjacent states. Empirical results on both public datasets and an online recommendation system demonstrate the effectiveness of the TADT-CSA model and its superiority over baseline methods.
@@ Introduction
Figure 1: Overall framework of the TADT-CSA model. $R$ denotes the temporal advantage signal, RP represents reward prediction, and CTP stands for contrastive transition prediction.
@@
From the RL perspective, the sequential recommendation problem can be modeled using a MDP. Formally, let $\mathcal{U}=\left\{u_{1},\cdots,u_{n}\right\}$ denote the set of users, and =$\mathcal{I}=\left\{i_{1},\cdots,i_{m}\right\}$ denote the candidate item set. The MDP for sequential recommendation can then be represented by the tuple $\left(\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\gamma\right)$, defined as follows:
@@
  • Action Space $\mathcal{A}$: The action $a_{t}\in\mathcal{A}$ corresponds to the recommended item list at time step $t$. For simplicity, we consider a recommendation list of length 1.
@@
[…]
Finally, the trajactory token representation list, i.e. $\left(\mathbf{R}_{1},\mathbf{c}_{1},\mathbf{a}_{1}\cdots,\mathbf{R}_{T},\mathbf{c}_{T},\mathbf{a}_{T}\right)$ are fed into a Causal Transformer model (Vaswani et al. 2017), such as GPT (Radford et al. 2018). The predicted action probability $p(a_{t}|\tau_{0:t-1},s_{t},R_{t})$ is formulated as follows:
$\hat{\mathbf{a}}_{t}=\text{GPT}(\mathbf{R}_{<t},\mathbf{c}_{<t},\mathbf{a}_{<t},\mathbf{R}_{t},\mathbf{c}_{t})$ (5)
@@
$\hat{p}(a_{t}\mid\tau_{0:t-1},s_{t},R_{t})=\text{softmax}_{\mathcal{I}}(\ell_{i})$
where $\mathcal{I}$ denotes the candidate item set, and $\mathbf{\Phi}(\cdot)$ is an item-kernel function that generates item embeddings. These components are not considered part of our model architecture.
@@
$\hat{\mathbf{R}}_{t}=\text{GPT}(\mathbf{R}_{<t},\mathbf{c}_{<t},\mathbf{a}_{<t})$ (6)
@@
To reduce the computational cost, we instead use the quantized codebook indices $c_{t}$ as compact representations of the original states $s_{t}$. Specifically, we apply a sequence hashing technique based on the transformed sub-trajectory $(c_{1},a_{1},R_{1},\cdots,c_{t})$, which significantly improves efficiency. For each batch, we hash these sub-trajectories into a set of groups:
@@
We then introduce a quantile-based pairwise ranking loss, designed to enhance policy learning within each group. For each group $g_{t,k}$, we compute the $\beta$-quantile of the RTG values and divide the samples into positive and negative subsets. The threshold is defined as $\text{threshold}_{\beta}=\text{Quantile}(R^{\text{RTG}}(g_{t,k}),\beta)$. To further improve computational efficiency, we adopt a top-K selection strategy in place of computing explicit quantile thresholds. Specifically, for each group $g_{t,k}$ of size $N_{g}$, we select the top $\lfloor\beta N_{g}\rfloor$ samples with the highest RTG values and treat them as positive examples. This avoids full sorting and reduces the time complexity from $O(n\log n)$ to $O(n)$.
@@
For the CTP network, we adopt a contrastive learning approach inspired by SimCLR (Chen et al. 2020), aiming to capture transition dynamics in the compressed space. Specifically, we treat the triplet $(\mathbf{c}_{t},a_{t},\mathbf{c}_{t+1}$) as a positive sample, while treating $(\mathbf{c}_{t},a_{t},\mathbf{c}^{\prime}_{t+1}$) as negative samples, where $\mathbf{c}^{\prime}_{t+1}$ denotes a randomly selected embedding satisfying $c^{\prime}_{t+1}\neq c_{t+1}$. In practice, we sample $K_{\text{neg}}$ negative examples for each positive pair during training.
@@
$\varepsilon_{r}=\max_{s,a}\left|r(s,a)-r_{g}(f_{\theta}(s),a)\right|,$ (19)
@@
$\varepsilon_{\mathcal{T}}=\max_{s,a,s^{\prime}}\left|\mathcal{T}(s,a,s^{\prime})-\mathcal{T}{h}(f{\theta}(s),a,f_{\theta}(s^{\prime}))\right|.$ (20)
@@
$V_{\pi^{*}}(s)-V_{\pi_{\Theta}}(s)\leq\frac{2}{(1-\gamma)^{2}}\left(\varepsilon_{r}+\kappa I^{\frac{d+2}{2d}}|\mathcal{C}|^{-\frac{1}{d}}+\frac{\gamma\varepsilon_{\mathcal{T}}|\mathcal{C}|}{1-\gamma}\right)$ (21)
[…]
This theorem guarantees that the proposed CSA module preserves MDP equivalence in the compressed state space with error independent of the original state space size $|\mathcal{S}|$. The bound depends only on: 1) fixed codebook size $|\mathcal{C}|$, 2) prediction errors of RP/CTP networks, and 3) the concentration of state embeddings. This explains CSA’s robustness in industrial scenarios where $|\mathcal{S}|$ is often extremely large and grows dynamically. Furthermore, this theoretical guarantee is highly relevant for DT, as they implicitly learn policies by reconstructing optimal behaviors conditioned on future returns. The RP and CTP networks preserve critical reward and transition structures, enabling DT to achieve near-optimal decision-making in the abstracted state space.
@@
The compared methods includes traditional offline RL methods, like CQL (Kumar et al. 2020), IQL (Kostrikov, Nair, and Levine 2021); sequential recommendation model, like SASRec (Kang and McAuley 2018) and BERT4Rec (Sun et al. 2019) and DT-based recommendation methods, like DT4Rec (Zhao et al. 2023), CDT4Rec (Wang et al. 2023a) and DT4IER (Liu et al. 2024b).
@@
We evaluate the parameter sensitivity of TADT-CSA on the KuaiRand dataset and find that a codebook size of 646464 and $\delta=0.3$ yield the best performance, as shown in Figure 3. Additional parameter sensitivity analyses are provided in the Appendix.
@@ Appendix A Appendix
[…]
@@
$|\displaystyle|r(s_{t},a_{t})-\mathcal{R}_{\mathcal{C}}(c_{t},a_{t})|$$\leq|r(s_{t},a_{t})-\hat{r}_{t}|+|\hat{r}_{t}-\mathcal{R}_{\mathcal{C}}(c_{t},a_{t})|$ (25)
@@
$(b_{2})$ $\leq\sum_{c_{t+1}}\frac{\varepsilon_{\mathcal{T}}R_{max}}{1-\gamma}\leq\frac{\varepsilon_{\mathcal{T}}|\mathcal{C}|}{1-\gamma}$ (32)
@@
$(b)\leq\frac{\varepsilon_{\mathcal{T}}|\mathcal{C}|}{1-\gamma}+\kappa_{2}I^{\frac{d+2}{2d}}|\mathcal{C}|^{-\frac{1}{d}}$ (36)
@@
$\Delta Q$ $\leq\varepsilon_{r}+\kappa_{1}I^{\frac{d+2}{2d}}|\mathcal{C}|^{-\frac{1}{d}}+\gamma\left(\frac{\varepsilon_{\mathcal{T}}|\mathcal{C}|}{1-\gamma}+\kappa_{2}I^{\frac{d+2}{2d}}|\mathcal{C}|^{-\frac{1}{d}}\right)$ (37)
$\leq\varepsilon_{r}+(\kappa_{1}+\gamma\kappa_{2})I^{\frac{d+2}{2d}}|\mathcal{C}|^{-\frac{1}{d}}+\frac{\gamma\varepsilon_{\mathcal{T}}|\mathcal{C}|}{1-\gamma}$
@@
$\Delta Q\leq\varepsilon_{r}+\kappa I^{\frac{d+2}{2d}}|\mathcal{C}|^{-\frac{1}{d}}+\frac{\gamma\varepsilon_{\mathcal{T}}|\mathcal{C}|}{1-\gamma}$ (38)
@@
$V_{\pi^{*}}(s)-V_{\pi_{\Theta}}(s)$ $\leq\frac{2}{(1-\gamma)^{2}}\left(\varepsilon_{r}+\kappa I^{\frac{d+2}{2d}}|\mathcal{C}|^{-\frac{1}{d}}+\frac{\gamma\varepsilon_{\mathcal{T}}|\mathcal{C}|}{1-\gamma}\right)$ (39)
@@
[…]
@@
[…]

### v2

@@ Abstract
With the rapid advancement of Transformer-based Large Language Models (LLMs), generative recommendation has shown great potential in enhancing both the accuracy and semantic understanding of modern recommender systems. Compared to LLMs, the Decision Transformer (DT) is a lightweight generative model applied to sequential recommendation tasks. However, DT faces challenges in trajectory stitching, often producing suboptimal trajectories. Moreover, due to the high dimensionality of user states and the vast state space inherent in recommendation scenarios, DT can incur significant computational costs and struggle to learn effective state representations. To overcome these issues, we propose a novel Temporal Advantage Decision Transformer with Contrastive State Abstraction (TADT-CSA) model. Specifically, we combine the conventional Return-To-Go (RTG) signal with a novel temporal advantage (TA) signal that encourages the model to capture both long-term returns and their sequential trend. Furthermore, we integrate a contrastive state abstraction module into the DT framework to learn more effective and expressive state representations. Within this module, we introduce a TA–conditioned State Vector Quantization (TAC-SVQ) strategy, where the TA score guides the state codebooks to incorporate contextual token information. Additionally, a reward prediction network and a contrastive transition prediction (CTP) network are employed to ensure that the state codebook preserves both the reward information of the current state and the transition information between adjacent states. Empirical results on both public datasets and an online recommendation system demonstrate the effectiveness of the TADT-CSA model and its superiority over baseline methods.
@@ Introduction
Figure 1: Overall framework of the TADT-CSA model. $R$ denotes the return-conditioned signal shown in Eq 4, RP represents reward prediction, and CTP stands for contrastive transition prediction.
@@
From the RL perspective, the sequential recommendation problem can be modeled using a MDP. Formally, let $\mathcal{U}=\left\{u_{1},\cdots,u_{n}\right\}$ denote the set of users, and $\mathcal{I}=\left\{i_{1},\cdots,i_{m}\right\}$ denote the candidate item set. The MDP for sequential recommendation can then be represented by the tuple $\left(\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\gamma\right)$, defined as follows:
@@
  • Action Space $\mathcal{A}$: The action $a_{t}\in\mathcal{A}$ corresponds to the recommended item list at time step $t$.
@@
As illustrated in Fig 1, a trajectory $\tau$ of length $T$ is decomposed into a sequence $(R_{t},s_{t},a_{t})$ with total length $3T$, which is then fed into the Embedding Layer to obtain the corresponding state embedding $\mathbf{s}_{t}$, action embedding $\mathbf{a^{\prime}}_{t}$ and return embedding $\mathbf{R}^{\prime}_{t}=[\mathbf{R}^{\text{RTG}}_{t},\mathbf{R}^{\text{TA}}_{t}]$. Then through the return encoder and action encoder, which are commonly modeled by Multi-Layer Perceptrons(MLPs), we can obtain the return token embedding $\mathbf{R}_{t}$ and action token embedding $\mathbf{a}_{t}$. Given the high dimensionality and large magnitude of the state space, particularly in large-scale RS, we introduce the Contrastive State Abstraction (CSA) module to compress the high-dimensional state embeddings $\mathbf{s}_{t}$ into low-dimensional vector-quantized embeddings $\mathbf{c}_{t}$ with codebook id $c_{t}$.
Finally, the trajactory token representation list, i.e. $\left(\mathbf{R}_{1},\mathbf{c}_{1},\mathbf{a}_{1}\cdots,\mathbf{R}_{T},\mathbf{c}_{T},\mathbf{a}_{T}\right)$ are fed into a Causal Transformer model (Vaswani et al. 2017), such as GPT (Radford et al. 2018). The predicted action probability $p(a_{t}|\tau_{1:t-1},s_{t},R_{t})$ is formulated as follows:
$\hat{\mathbf{a}}_{t}=\text{GPT}(\mathbf{R}_{1:t-1},\mathbf{c}_{1:t-1},\mathbf{a}_{1:t-1},\mathbf{R}_{t},\mathbf{c}_{t})$ (5)
@@
$\hat{p}(a_{t}\mid\tau_{1:t-1},s_{t},R_{t})=\text{softmax}_{\mathcal{I}}(\ell_{i})$
where $\mathcal{I}$ denotes the candidate item set, and $\mathbf{\Phi}(\cdot)$ is a learnable item embedding layer that generates item embeddings.
@@
$\hat{\mathbf{R}}_{t}=\text{GPT}(\mathbf{R}_{1:t-1},\mathbf{c}_{1:t-1},\mathbf{a}_{1:t-1})$ (6)
@@
To reduce the computational cost of sub-trajectory matching, we instead use the quantized codebook indices $c_{t}$ as compact representations of the original states $s_{t}$. Specifically, we apply a sequence hashing technique based on the transformed sub-trajectory $(c_{1},a_{1},R_{1},\cdots,c_{t})$, which significantly improves efficiency. For each batch, we hash these sub-trajectories into a set of groups:
@@
We then introduce a quantile-based pairwise ranking loss, designed to enhance policy learning within each group. For each group $g_{t,k}$, we compute the $\beta$-quantile of the RTG values and divide the samples into positive and negative subsets. The threshold is defined as $\text{threshold}_{\beta}=\text{Quantile}(R^{\text{RTG}}(g_{t,k}),\beta)$. To further improve computational efficiency, we adopt a top-K selection strategy in place of computing explicit quantile thresholds. Specifically, for each group $g_{t,k}$ of size $N_{g}$, we adopt a QuickSelect algorithm, which is a quicksort-like partition-based method, to efficiently select the top $\lfloor\beta N_{g}\rfloor$ samples with the highest RTG values and treat them as positive examples. This avoids full sorting and reduces the time complexity from $O(n\log n)$ to $O(n)$.
@@
For the CTP network, we adopt a contrastive learning approach inspired by SimCLR (Chen et al. 2020), aiming to capture transition dynamics in the compressed space. Specifically, we treat the triplet $(\mathbf{c}_{t},a_{t},\mathbf{c}_{t+1}$) as a positive sample, while treating $(\mathbf{c}_{t},a_{t},\mathbf{c}^{\prime}_{t+1}$) as negative samples, where $\mathbf{c}^{\prime}_{t+1}$ denotes a randomly selected embedding satisfying $c^{\prime}_{t+1}\neq c_{t+1}$ and $c^{\prime}_{t+1}\neq c_{t}$. In practice, we sample $K_{\text{neg}}$ negative examples for each positive pair during training.
@@
$\varepsilon_{r}=\max_{s,a}\left|r(s,a)-r_{g_{\phi}}(f_{\theta}(s),a)\right|,$ (19)
@@
$\varepsilon_{\mathcal{P}}=\max_{s,a,s^{\prime}}\left|\mathcal{P}(s,a,s^{\prime})-\mathcal{P}_{h_{\psi}}(f_{\theta}(s),a,f_{\theta}(s^{\prime}))\right|.$ (20)
@@
$V_{\pi^{*}}(s)-V_{\pi_{\Theta}}(s)\leq\frac{2}{(1-\gamma)^{2}}\left(\varepsilon_{r}+\kappa I^{\frac{d+2}{2d}}|\mathcal{C}|^{-\frac{1}{d}}+\frac{\gamma\varepsilon_{\mathcal{P}}|\mathcal{C}|}{1-\gamma}\right)$ (21)
where $\kappa$ is a Lipschitz constant, $I=\int p(\mathbf{e})^{\frac{d}{d+2}}d\mathbf{e}$ is the distribution concentration factor ($0<I\leq 1$), $d$ is the embedding dimension and $|\mathcal{C}|$ is the codebook size.
This theorem guarantees that the proposed CSA module preserves MDP equivalence in the compressed state space, with approximation error independent of original state space size $|\mathcal{S}|$. The error bound depends only on: (1) the fixed codebook size $|\mathcal{C}|$, (2) the prediction errors of the RP and CTP networks, and (3) the concentration of the state embeddings. The property explains the robustness of CSA in industrial recommendation scenarios, where the state space is typically extremely large and dynamically evolving. Furthermore, this theoretical guarantee is particularly well-aligned with the inductive bias of DTs, which implicitly learn policies by reconstructing high-return trajectories conditioned on future RTG signals. By preserving critical reward and transition dynamics through the RP and CTP networks, the abstracted state space retains sufficient MDP structure, thereby enabling DTs to achieve near-optimal decision-making even under compressed representations.
@@
The compared methods include traditional offline RL methods, like CQL (Kumar et al. 2020), IQL (Kostrikov, Nair, and Levine 2021); sequential recommendation model, like SASRec (Kang and McAuley 2018) and BERT4Rec (Sun et al. 2019) and DT-based recommendation methods, like DT4Rec (Zhao et al. 2023), CDT4Rec (Wang et al. 2023a) and DT4IER (Liu et al. 2024b).
@@
We evaluate the parameter sensitivity of TADT-CSA on the KuaiRand dataset and find that a codebook size of $64$ and $\delta=0.3$ yield the best performance, as shown in Figure 3. Additional parameter sensitivity analyses are provided in the Appendix.
@@ Appendix A Appendix
Suppose the state embedding distribution p(e) has compact support and satisfies $p(\mathbf{e})\geq\rho_{\min}>0$ on its support. Define $I=\int p(\mathbf{e})^{\frac{d}{d+2}}d\mathbf{e}$. Then there exists a codebook $C$ with $|C|=N$ such that the covering radius satisfies:
@@
$|r(s_{t},a_{t})-\mathcal{R}_{\mathcal{C}}(c_{t},a_{t})|$ $\leq|r(s_{t},a_{t})-\hat{r}_{t}|+|\hat{r}_{t}-\mathcal{R}_{\mathcal{C}}(c_{t},a_{t})|\$(25)
@@
$(b_{2})$ $\leq\sum_{c_{t+1}}\frac{\varepsilon_{\mathcal{P}}R_{max}}{1-\gamma}\leq\frac{\varepsilon_{\mathcal{P}}|\mathcal{C}|}{1-\gamma}$ (32)
@@
$(b)\leq\frac{\varepsilon_{\mathcal{P}}|\mathcal{C}|}{1-\gamma}+\kappa_{2}I^{\frac{d+2}{2d}}|\mathcal{C}|^{-\frac{1}{d}}$ (36)
@@
$\Delta Q$ $\leq\varepsilon_{r}+\kappa_{1}I^{\frac{d+2}{2d}}|\mathcal{C}|^{-\frac{1}{d}}+\gamma\left(\frac{\varepsilon_{\mathcal{P}}|\mathcal{C}|}{1-\gamma}+\kappa_{2}I^{\frac{d+2}{2d}}|\mathcal{C}|^{-\frac{1}{d}}\right)$ (37)
$\leq\varepsilon_{r}+(\kappa_{1}+\gamma\kappa_{2})I^{\frac{d+2}{2d}}|\mathcal{C}|^{-\frac{1}{d}}+\frac{\gamma\varepsilon_{\mathcal{P}}|\mathcal{C}|}{1-\gamma}$
@@
$\Delta Q\leq\varepsilon_{r}+\kappa I^{\frac{d+2}{2d}}|\mathcal{C}|^{-\frac{1}{d}}+\frac{\gamma\varepsilon_{\mathcal{P}}|\mathcal{C}|}{1-\gamma}$ (38)
@@
$V_{\pi^{*}}(s)-V_{\pi_{\Theta}}(s)$ $\leq\frac{2}{(1-\gamma)^{2}}\left(\varepsilon_{r}+\kappa I^{\frac{d+2}{2d}}|\mathcal{C}|^{-\frac{1}{d}}+\frac{\gamma\varepsilon_{\mathcal{P}}|\mathcal{C}|}{1-\gamma}\right)$ (39)
@@
    $\text{Recall@}K=\frac{|\{\text{relevant items}\}\cap\{\text{top-K recommendations}\}|}{|\{\text{relevant items}\}|}$
@@
    $\text{NDCG@}K=\frac{\text{DCG@}K}{\text{IDCG@}K},\quad\text{DCG@}K=\sum_{i=1}^{K}\frac{2^{rel_{i}}-1}{\log_{2}(i+1)}$
@@
Computational Cost Analysis
We compared the computational costs of different methods on the KuaiRand-Pure training dataset, as shown in Table 4. The results indicate that DT-based methods are more computationally efficient than LLM-based methods, highlighting the lightweight nature of DT models in terms of shorter running time, lower GPU memory consumption, and faster convergence. Specifically, the running time and GPU memory usage of TADT-CSA are comparable to those of existing DT-based methods, demonstrating that the inclusion of the CSA module does not increase model complexity or computational overhead.
Models #Running Time #GPU Memory #Convergence Steps
DT4Rec 4.5h 8GB 14,280
DT4IER 3.9h 8GB 12,380
CDT4Rec 3.4h 8GB 9,600
LLM 18.6h 24GB 307,200
TADT-CSA 4.0h 12GB 8,600
Table 4: Computational cost of different methods.
Figure 5: Trajectory stitching evaluation of different DT-based recommendation methods.
Trajectory Stitching Evaluation
Since Decision Transformer (DT) models often suffer from the trajectory stitching problem, as discussed in prior works such as ACT (Gao et al. 2024), we conduct experiments on the KuaiRand-Pure dataset to evaluate how different chunk lengths impact model performance. Specifically, we divide full-length trajectories (length = 200) into smaller chunks of fixed lengths (e.g., 30, 50) and compare the cumulative reward achieved by TADT-CSA against other DT-based baselines, including DT4Rec, CDT4Rec and DT4IER.
As shown in Figure 5, TADT-CSA consistently outperforms all baselines across different chunk sizes, and maintains strong performance even when using the full trajectories. In contrast, other DT-based methods exhibit noticeably degraded performance when using full-length trajectories, despite the fact that full sequences are typically expected to yield richer information and better results. We hypothesize that this counterintuitive behavior stems from the fact that excessively long sequences (e.g., 200 steps) pose significant optimization and modeling challenges for standard DTs, such as over-smoothing or failure to attend to earlier tokens effectively. This makes them more vulnerable to trajectory stitching issues in the full-length setting.
Moreover, DT4Rec and CDT4Rec perform relatively well on shorter chunks (e.g., 30), but their performance drops and convergence slows as chunk size increases. DT4IER, on the other hand, demonstrates stable convergence across all chunk sizes (30, 50, and full), with its cumulative reward reaching similar levels in each case. However, it still underperforms compared to TADT-CSA, indicating that despite its robustness, DT4IER remains slightly less effective in capturing long-term value signals.
These findings demonstrate the implicit trajectory stitching capability of TADT-CSA and its superior ability to model long-horizon dependencies in sequential recommendation scenarios.